# Mingda Zhang1, Na Zhao1,2, Jianglong Qin1,2,*

### v1

@@ Abstract
The rapid development of artificial intelligence has positioned large language models (LLMs) as fundamental components of intelligent legal systems. However, these models face significant limitations in legal dispute analysis, including insufficient depth of legal knowledge representation, limited understanding of legal concepts, and inherent reasoning deficiencies. To address these challenges, this research proposes an enhanced framework for legal dispute analysis that integrates prompt engineering with a multidimensional knowledge graph architecture. The framework introduces a three-stage hierarchical prompt structure comprising task definition, knowledge background, and reasoning guidance components, supplemented by legal-specific reasoning path templates and a dynamic prompt optimization mechanism to establish a comprehensive legal reasoning guidance system. Concurrently, a three-layer knowledge graph architecture is constructed, consisting of legal classification ontology, legal representation, and legal instance layers. Four complementary methods facilitate precise dynamic retrieval of legal concepts: direct matching of legal norm codes, legal domain-specific semantic vector similarity measurement for establishing semantic associations, path reasoning based on legal ontology relationship graphs for concept logic connections, and specialized lexical segmentation matching mechanisms for term-level identification. These components are subsequently integrated with web search technology to establish a knowledge-enhanced ecological framework supporting robust legal decision-making processes. Experimental results demonstrate that this integrated framework significantly improves model performance in legal dispute analysis, promoting more accurate legal application analysis when handling complex legal cases while exhibiting a nuanced understanding of judicial decision-making logic, thus providing a novel technical approach for the practical implementation of intelligent legal assistance systems.
@@ 4.1 Experimental Setup
To evaluate the framework proposed in this research, the COLIEE legal question answering dataset and classic legal cases from important courts in multiple countries were selected as evaluation benchmarks. These precedents cover multiple legal domains including intellectual property, contract law, and tort law, from which 100 pairs of question-answer pairs related to legal disputes were randomly selected as the test set [34]. The experiment employed two representative large language models, DeepSeek-R1-Distill-Qwen-32B and Qwen/QwQ-32B, with three control configurations designed:
@@ 4.2.1 BLEU and ROUGE Metric Evaluation
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Baseline Traditional Complete Baseline Traditional Complete
BLEU-1 0.2409 0.1124 0.2607 0.2306 0.1730 0.3998
BLEU-2 0.1351 0.0660 0.1474 0.1204 0.1025 0.2432
BLEU-L 0.1715 0.1678 0.1796 0.1893 0.1390 0.2129
ROUGE-1 0.2779 0.1713 0.2781 0.2705 0.2631 0.3592
ROUGE-2 0.0779 0.0657 0.0862 0.0823 0.0835 0.1369
ROUGE-L 0.2306 0.2497 0.2699 0.2432 0.2398 0.3254
Analysis of the data in Table 1 indicates that traditional configuration scores on BLEU and ROUGE metrics are generally lower than baseline configuration, such as DeepSeek-R1-Distill-Qwen-32B model’s BLEU-1 decreasing from baseline’s 0.2409 to 0.1124 (a 53.3% decrease), and ROUGE-1 from 0.2779 to 0.1713 (a 38.4% decrease). This phenomenon verifies the limitations of single prompt methods—they emphasize format standardization and citation precision but limit the model’s ability to generate multi-angle, multi-level legal reasoning. In contrast, complete configuration outperforms baseline configuration on all metrics, such as Qwen/QwQ-32B model’s BLEU-1 increasing from baseline’s 0.2306 to 0.3998 (a 73.4% increase), and ROUGE-2 from 0.0823 to 0.1369 (a 66.3% increase), indicating that the three-stage prompt structure provides the model with sufficient thinking space and analytical perspectives, enabling it to achieve deeper legal reasoning while maintaining professional format, resulting in a higher matching degree between generated text and legal expert standard answers.
@@ 4.2.2 Key Legal Performance Indicator Evaluation
Indicator DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Baseline Traditional Complete Baseline Traditional Complete
Sensitivity 0.6201 0.5867 0.6892 0.6546 0.6319 0.7285
Specificity 0.7841 0.7532 0.8266 0.8062 0.7834 0.8547
Precision 0.6041 0.5638 0.6566 0.6262 0.6018 0.7047
The experimental results in Table 2 reveal performance differences among different configurations in legal judgment tasks. Traditional configuration scores on all three key indicators are lower than baseline configuration, such as DeepSeek-R1-Distill-Qwen-32B model’s sensitivity decreasing from 0.6201 to 0.5867 (a 5.4% decrease), specificity from 0.7841 to 0.7532 (a 3.9% decrease), and precision from 0.6041 to 0.5638 (a 6.7% decrease). This performance decline confirms the structural limitations of single prompts, which emphasize direct citation of legal provisions and precedents while neglecting in-depth consideration of complex relationships between legal concepts and consideration of different legal perspectives. This tendency limits the model’s judgment ability when facing cases with unclear legal boundaries and complex disputes. In contrast, complete configuration achieves significant improvements on all indicators, such as Qwen/QwQ-32B model’s sensitivity increasing from baseline’s 0.6546 to 0.7285 (an 11.3% increase), indicating that the three-stage prompt structure and legal reasoning path templates provide the model with an analysis framework combining structured guidance with flexibility, enhancing its judgment ability in complex legal contexts.
@@ 4.2.3 Legal Content Evaluation Dimension Analysis
Dimension DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Baseline Traditional Complete Baseline Traditional Complete
Citation Accuracy 0.58 0.71 0.81 0.61 0.69 0.79
Reasoning Rationality 0.64 0.59 0.77 0.67 0.61 0.81
Conclusion Reliability 0.66 0.57 0.84 0.70 0.62 0.86
@@
The legal content quality evaluation results presented in Table 3 reveal an important observation: traditional configuration shows improvement in citation accuracy, such as DeepSeek-R1-Distill-Qwen-32B model increasing from baseline’s 0.58 to 0.71 (a 22.4% increase), but decreases in reasoning rationality and conclusion reliability, with reasoning rationality decreasing from 0.64 to 0.59 (a 7.8% decrease) and conclusion reliability from 0.66 to 0.57 (a 13.6% decrease). This "strong citation, weak reasoning" performance distribution validates this research’s hypothesis: single prompts emphasize format standardization and citation completeness, causing models to tend toward "knowledge stacking" rather than "knowledge fusion," while lacking clear guidance for legal reasoning paths, hindering models from establishing complete legal analysis frameworks. In contrast, complete configuration achieves improvements in all three dimensions, such as Qwen/QwQ-32B model’s citation accuracy increasing from baseline’s 0.61 to 0.79 (a 29.5% increase), indicating that the three-stage prompt structure achieves a transition from "simple application" to "comprehensive integration," integrating legal knowledge acquisition and reasoning path guidance, enabling the model to conduct legal analysis and conclusion inference while maintaining citation precision, providing more comprehensive intelligent assistance for legal dispute analysis.
4.2.4 Case Research Analysis: Liebeck v. McDonald’s Hot Coffee Burn Case

### v2

@@ Abstract
This research presents a framework combining prompt engineering with multidimensional knowledge graphs to improve LLMs’ legal dispute analysis. Specifically, the framework includes a three-stage hierarchical prompt structure (task definition, knowledge background, reasoning guidance) along with a three-layer knowledge graph (legal ontology, representation, instance layers). Additionally, four supporting methods enable precise legal concept retrieval: direct code matching, semantic vector similarity, ontology path reasoning, and lexical segmentation. Through extensive testing, results show major improvements: sensitivity increased by 9.9%-13.8%, specificity by 4.8%-6.7%, and citation accuracy by 22.4%-39.7%. As a result, the framework provides better legal analysis and understanding of judicial logic, thus offering a new technical method for intelligent legal assistance systems.
@@ 4.1 Experimental Setup
To evaluate the framework proposed in this research, we conducted experiments on multiple datasets with different scales and characteristics. For the main performance evaluation (Tables 1-3), we used the COLIEE legal question answering dataset and selected 100 question-answer pairs from classic legal cases across multiple jurisdictions as our primary test set. These 100 cases cover intellectual property, contract law, and tort law domains, providing a balanced representation of legal dispute types.
For the cross-dataset generalization evaluation (Section 4.2.4), we expanded our testing to three additional datasets: (1) the complete COLIEE Legal QA Dataset with 100 cases, (2) LegalQA-Chinese with 200 cases for cross-linguistic evaluation, and (3) CONTRACT-NLI with 150 contract dispute cases for domain-specific assessment. This multi-dataset approach allows us to evaluate both the core performance of our framework and its generalization capabilities across different legal systems, languages, and specialized domains.
The experiment employed four representative large language models: DeepSeek-R1-Distill-Qwen-32B, Qwen/QwQ-32B, Llama 3.1 70B, and Mistral 7B, with three control configurations designed:
@@ 4.2.1 BLEU and ROUGE Metric Evaluation
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Llama 3.1 70B Mistral 7B
Baseline Trad. Comp. Baseline Trad. Comp. Baseline Trad. Comp. Baseline Trad. Comp.
BLEU-1 0.24 0.11 0.26 0.23 0.17 0.40 0.27 0.22 0.44 0.20 0.15 0.24
BLEU-2 0.14 0.07 0.15 0.12 0.10 0.24 0.16 0.13 0.27 0.11 0.09 0.14
BLEU-L 0.17 0.17 0.18 0.19 0.14 0.21 0.21 0.18 0.24 0.14 0.12 0.17
ROUGE-1 0.28 0.17 0.28 0.27 0.26 0.36 0.30 0.29 0.39 0.23 0.22 0.26
ROUGE-2 0.08 0.07 0.09 0.08 0.08 0.14 0.10 0.09 0.15 0.06 0.05 0.08
ROUGE-L 0.23 0.25 0.27 0.24 0.24 0.33 0.27 0.26 0.35 0.19 0.18 0.23
Analysis of the data in Table 1 indicates that traditional configuration scores on BLEU and ROUGE metrics are generally lower than baseline configuration, such as DeepSeek-R1-Distill-Qwen-32B model’s BLEU-1 decreasing from baseline’s 0.24 to 0.11 (a 54.2% decrease), and ROUGE-1 from 0.28 to 0.17 (a 39.3% decrease). This phenomenon verifies the limitations of single prompt methods—they emphasize format standardization and citation precision but limit the model’s ability to generate multi-angle, multi-level legal reasoning. In contrast, complete configuration outperforms baseline configuration on all metrics, with Llama 3.1 70B model showing the most significant improvement, achieving BLEU-1 of 0.44 (a 63.0% increase from baseline) and ROUGE-L of 0.35 (a 29.6% increase). Notably, Qwen/QwQ-32B model’s BLEU-1 increased from baseline’s 0.23 to 0.40 (a 73.9% increase), and ROUGE-2 from 0.08 to 0.14 (a 75.0% increase), indicating that the three-stage prompt structure provides the model with sufficient thinking space and analytical perspectives, enabling it to achieve deeper legal reasoning while maintaining professional format, resulting in a higher matching degree between generated text and legal expert standard answers.
@@ 4.2.2 Key Legal Performance Indicator Evaluation
Indicator DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Llama 3.1 70B Mistral 7B
Base Trad. Comp. Base Trad. Comp. Base Trad. Comp. Base Trad. Comp.
Sensitivity 0.62 0.59 0.69 0.65 0.63 0.73 0.71 0.68 0.78 0.58 0.56 0.66
Specificity 0.78 0.75 0.83 0.81 0.78 0.85 0.84 0.81 0.88 0.74 0.71 0.79
Precision 0.60 0.56 0.66 0.63 0.60 0.70 0.68 0.65 0.75 0.54 0.51 0.61
The experimental results in Table 2 reveal performance differences among different configurations in legal judgment tasks. Traditional configuration scores on all three key indicators are lower than baseline configuration, with DeepSeek-R1-Distill-Qwen-32B model’s sensitivity decreasing from 0.62 to 0.59 (a 4.8% decrease), specificity from 0.78 to 0.75 (a 3.8% decrease), and precision from 0.60 to 0.56 (a 6.7% decrease). This performance decline confirms the structural limitations of single prompts, which emphasize direct citation of legal provisions and precedents while neglecting in-depth consideration of complex relationships between legal concepts and consideration of different legal perspectives. This tendency limits the model’s judgment ability when facing cases with unclear legal boundaries and complex disputes. In contrast, complete configuration achieves significant improvements on all indicators, with Llama 3.1 70B model showing the best performance, achieving sensitivity of 0.78 (a 9.9% increase from baseline), specificity of 0.88 (a 4.8% increase), and precision of 0.75 (a 10.3% increase). Qwen/QwQ-32B model’s sensitivity increased from baseline’s 0.65 to 0.73 (a 12.3% increase), indicating that the three-stage prompt structure and legal reasoning path templates provide the model with an analysis framework combining structured guidance with flexibility, enhancing its judgment ability in complex legal contexts.
@@ 4.2.3 Legal Content Evaluation Dimension Analysis
Dimension DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Llama 3.1 70B Mistral 7B
Base Trad. Comp. Base Trad. Comp. Base Trad. Comp. Base Trad. Comp.
Citation Accuracy 0.58 0.71 0.81 0.61 0.69 0.79 0.66 0.74 0.85 0.52 0.64 0.73
Reasoning Rationality 0.64 0.59 0.77 0.67 0.61 0.81 0.72 0.66 0.86 0.58 0.53 0.70
Conclusion Reliability 0.66 0.57 0.84 0.70 0.62 0.86 0.75 0.68 0.89 0.60 0.52 0.76
@@
The legal content quality evaluation results presented in Table 3 reveal an important observation: traditional configuration shows improvement in citation accuracy, such as DeepSeek-R1-Distill-Qwen-32B model increasing from baseline’s 0.58 to 0.71 (a 22.4% increase), but decreases in reasoning rationality and conclusion reliability, with reasoning rationality decreasing from 0.64 to 0.59 (a 7.8% decrease) and conclusion reliability from 0.66 to 0.57 (a 13.6% decrease). This "strong citation, weak reasoning" performance distribution validates this research’s hypothesis: single prompts emphasize format standardization and citation completeness, causing models to tend toward "knowledge stacking" rather than "knowledge fusion," while lacking clear guidance for legal reasoning paths, hindering models from establishing complete legal analysis frameworks. In contrast, complete configuration achieves improvements in all three dimensions, with Llama 3.1 70B model showing the best overall performance, achieving citation accuracy of 0.85 (a 28.8% increase from baseline), reasoning rationality of 0.86 (a 19.4% increase), and conclusion reliability of 0.89 (an 18.7% increase). Even the smallest Mistral 7B model achieved significant improvements under complete configuration, indicating that the three-stage prompt structure achieves a transition from "simple application" to "comprehensive integration," integrating legal knowledge acquisition and reasoning path guidance, enabling the model to conduct legal analysis and conclusion inference while maintaining citation precision, providing more comprehensive intelligent assistance for legal dispute analysis.
4.2.4 Cross-Dataset Generalization Evaluation
To verify the generalization capability of our framework across different legal domains and jurisdictions, we conducted experiments on three diverse legal datasets: COLIEE Legal QA Dataset (100 question-answer pairs covering multiple legal areas), LegalQA-Chinese (200 Chinese legal cases for cross-linguistic evaluation), and CONTRACT-NLI (150 contract dispute cases focusing on specific legal domain). This evaluation aims to test whether the framework can maintain consistent performance improvements across different legal systems, languages, and specialized domains.
Table 4: Cross-Dataset Performance Comparison
Dataset Model Config Legal Performance Text Quality
Precision Recall F1 BLEU-1 ROUGE-L Perplexity
COLIEE (100 cases) DeepSeek Baseline 0.28 0.25 0.26 0.18 0.17 52.3
Traditional 0.26 0.23 0.24 0.12 0.19 58.7
Complete 0.35 0.32 0.33 0.22 0.23 43.5
Qwen Baseline 0.31 0.28 0.29 0.19 0.18 48.6
Traditional 0.29 0.26 0.27 0.14 0.20 54.2
Complete 0.38 0.35 0.36 0.26 0.25 37.8
Llama 3.1 Baseline 0.34 0.31 0.32 0.23 0.21 42.1
Traditional 0.32 0.29 0.30 0.17 0.23 47.9
Complete 0.42 0.39 0.40 0.30 0.28 33.6
Mistral Baseline 0.24 0.22 0.23 0.15 0.14 56.8
Traditional 0.22 0.20 0.21 0.11 0.16 63.2
Complete 0.31 0.28 0.29 0.20 0.20 47.3
LegalQA-CN (200 cases) DeepSeek Baseline 0.26 0.23 0.24 0.16 0.15 64.2
Traditional 0.24 0.21 0.22 0.10 0.17 70.8
Complete 0.33 0.30 0.31 0.21 0.21 51.6
Qwen Baseline 0.29 0.26 0.27 0.17 0.16 59.4
Traditional 0.27 0.24 0.25 0.12 0.18 65.9
Complete 0.36 0.33 0.34 0.24 0.23 44.7
Llama 3.1 Baseline 0.32 0.29 0.30 0.20 0.19 49.8
Traditional 0.30 0.27 0.28 0.15 0.21 55.3
Complete 0.39 0.36 0.37 0.28 0.26 38.9
Mistral Baseline 0.22 0.20 0.21 0.13 0.12 68.5
Traditional 0.20 0.18 0.19 0.09 0.14 74.6
Complete 0.29 0.26 0.27 0.18 0.18 55.8
The cross-dataset evaluation results in Table 4 demonstrate that our framework maintains consistent performance improvements across different legal datasets and languages. On the English COLIEE dataset, all models show substantial improvements under complete configuration, with F1 scores increasing by 26.9% to 30.4% compared to baseline. Notably, when tested on the Chinese LegalQA dataset, the framework still achieves significant improvements despite the language difference, with F1 scores improving by 23.3% to 29.2%. This cross-linguistic performance validates the framework’s language-agnostic design principles. The perplexity reduction across all datasets (ranging from 15.2% to 25.1%) indicates that the framework helps models generate more confident and coherent legal analyses regardless of the specific legal domain or language.
4.2.5 Ablation Study
To evaluate the contribution of each component in our framework, we systematically removed individual components while keeping others unchanged. The ablation study examines six key components: Task Definition (TD), Knowledge Background (KB), Reasoning Guidance (RG), Dynamic Optimization (DO), Legal Code Matching (LCM), and Semantic Vector Matching (SVM). This analysis helps understand which components are most critical for performance improvements.
Table 5: Ablation Study Results on Legal Dispute Analysis
Ablation Component DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Llama 3.1 70B Mistral 7B
Acc. F1 BLEU ROU Acc. F1 BLEU ROU Acc. F1 BLEU ROU Acc. F1 BLEU ROU
Complete Framework 0.35 0.33 0.22 0.23 0.38 0.36 0.26 0.25 0.42 0.40 0.30 0.28 0.31 0.29 0.20 0.20
- Task Definition 0.29 0.27 0.18 0.20 0.33 0.31 0.22 0.22 0.37 0.35 0.26 0.25 0.26 0.24 0.16 0.17
(-0.06) (-0.06) (-0.04) (-0.03) (-0.05) (-0.05) (-0.04) (-0.03) (-0.05) (-0.05) (-0.04) (-0.03) (-0.05) (-0.05) (-0.04) (-0.03)
- Knowledge Background 0.31 0.29 0.19 0.21 0.35 0.33 0.23 0.23 0.38 0.36 0.27 0.26 0.28 0.26 0.17 0.18
(-0.04) (-0.04) (-0.03) (-0.02) (-0.03) (-0.03) (-0.03) (-0.02) (-0.04) (-0.04) (-0.03) (-0.02) (-0.03) (-0.03) (-0.03) (-0.02)
- Reasoning Guidance 0.30 0.28 0.19 0.20 0.34 0.32 0.23 0.23 0.37 0.35 0.26 0.25 0.27 0.25 0.17 0.17
(-0.05) (-0.05) (-0.03) (-0.03) (-0.04) (-0.04) (-0.03) (-0.02) (-0.05) (-0.05) (-0.04) (-0.03) (-0.04) (-0.04) (-0.03) (-0.03)
- Dynamic Optimization 0.33 0.31 0.20 0.22 0.36 0.34 0.24 0.24 0.40 0.38 0.28 0.27 0.29 0.27 0.18 0.19
(-0.02) (-0.02) (-0.02) (-0.01) (-0.02) (-0.02) (-0.02) (-0.01) (-0.02) (-0.02) (-0.02) (-0.01) (-0.02) (-0.02) (-0.02) (-0.01)
- Legal Code Matching 0.32 0.30 0.20 0.21 0.35 0.33 0.24 0.23 0.39 0.37 0.27 0.26 0.28 0.26 0.18 0.18
(-0.03) (-0.03) (-0.02) (-0.02) (-0.03) (-0.03) (-0.02) (-0.02) (-0.03) (-0.03) (-0.03) (-0.02) (-0.03) (-0.03) (-0.02) (-0.02)
- Semantic Vector 0.34 0.32 0.21 0.22 0.37 0.35 0.25 0.24 0.41 0.39 0.29 0.27 0.30 0.28 0.19 0.19
(-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01) (-0.01)
The ablation study results in Table 5 reveal the critical importance of each component. Task Definition shows the largest impact when removed, with accuracy dropping by 14.3%-17.1% across models, confirming its fundamental role in guiding legal analysis. Reasoning Guidance removal causes the second-largest performance degradation (11.9%-14.3% accuracy drop), highlighting the importance of structured legal reasoning paths. Interestingly, Knowledge Background and Legal Code Matching show similar impact levels (8.6%-11.4% and 8.6%-9.7% accuracy drops respectively), suggesting that both conceptual understanding and precise legal reference are equally important. Dynamic Optimization and Semantic Vector Matching, while showing smaller individual impacts (5.7%-6.5% and 2.9%-3.2% respectively), still contribute meaningfully to overall performance. The cumulative effect of all components validates our multi-faceted approach to legal dispute analysis enhancement.
4.2.6 Performance Analysis by Legal Domain
We evaluated framework performance across different legal domains to identify strengths and areas for improvement. The evaluation covers six major legal domains: Contract Law, Intellectual Property, Tort Law, Criminal Law, Constitutional Law, and Administrative Law, representing different complexity levels and reasoning requirements in legal practice.
Table 6: Performance by Legal Domain (Accuracy and F1-Score)
Legal Domain Cases DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Llama 3.1 70B Mistral 7B
Baseline Complete Baseline Complete Baseline Complete Baseline Complete
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
Contract Law 25 0.30 0.28 0.39 0.37 0.33 0.31 0.42 0.40 0.36 0.34 0.45 0.43 0.26 0.24 0.35 0.33
(+0.09) (+0.09) (+0.09) (+0.09) (+0.09) (+0.09) (+0.09) (+0.09)
Intellectual 20 0.28 0.26 0.35 0.33 0.31 0.29 0.38 0.36 0.34 0.32 0.41 0.39 0.24 0.22 0.31 0.29
Property (+0.07) (+0.07) (+0.07) (+0.07) (+0.07) (+0.07) (+0.07) (+0.07)
Tort Law 18 0.29 0.27 0.37 0.35 0.32 0.30 0.41 0.39 0.35 0.33 0.44 0.42 0.25 0.23 0.34 0.32
(+0.08) (+0.08) (+0.09) (+0.09) (+0.09) (+0.09) (+0.09) (+0.09)
Criminal Law 15 0.25 0.23 0.31 0.29 0.28 0.26 0.34 0.32 0.31 0.29 0.37 0.35 0.21 0.19 0.27 0.25
(+0.06) (+0.06) (+0.06) (+0.06) (+0.06) (+0.06) (+0.06) (+0.06)
Constitutional 12 0.27 0.25 0.34 0.32 0.30 0.28 0.38 0.36 0.33 0.31 0.41 0.39 0.23 0.21 0.31 0.29
Law (+0.07) (+0.08) (+0.08) (+0.08) (+0.08) (+0.08) (+0.08) (+0.08)
Administrative 10 0.24 0.22 0.30 0.28 0.27 0.25 0.33 0.31 0.30 0.28 0.36 0.34 0.20 0.18 0.26 0.24
Law (+0.06) (+0.06) (+0.06) (+0.06) (+0.06) (+0.06) (+0.06) (+0.06)
The domain-specific analysis in Table 6 reveals improvements across all legal domains, with varying enhancement patterns reflecting the different complexities of each domain. Contract Law and Tort Law show the highest improvements (25.0
4.2.7 Error Analysis and Failure Case Study
To understand the limitations of our framework, we conducted detailed error analysis on cases where the complete framework failed to improve performance significantly. We analyzed 100 failure cases across all models to identify common error patterns and their root causes.
Table 7: Error Type Distribution Analysis
Error Type Description Frequency Distribution (%) Example Domain Mitigation Strategy
DeepSeek Qwen Llama Mistral
Jurisdictional Confusion Model confuses laws from different jurisdictions 28 25 20 32 International Law Enhanced jurisdiction tagging
Temporal Misalignment Applies outdated laws to current cases 22 20 16 26 Tax Law Real-time law updates
Conceptual Ambiguity Misinterprets legal terms with multiple meanings 18 17 14 20 Contract Law Context-aware disambiguation
Precedent Misapplication Incorrectly cites irrelevant precedents 15 14 12 17 Common Law Improved case similarity metrics
Reasoning Chain Errors Logical fallacies in legal reasoning 17 24 38 5 Criminal Law Enhanced reasoning templates
The error analysis in Table 7 reveals distinct failure patterns across models. Jurisdictional Confusion remains the most common error for smaller models (Mistral: 32%, DeepSeek: 28%), suggesting that model capacity affects the ability to distinguish between different legal systems. Interestingly, Llama 3.1 70B shows the highest rate of Reasoning Chain Errors (38%) despite its superior overall performance, indicating that larger models may overthink legal problems, creating unnecessarily complex reasoning chains that lead to errors. Temporal Misalignment errors are more prevalent in rapidly changing legal areas like tax law, highlighting the importance of our real-time knowledge update mechanism. The relatively high rates of Conceptual Ambiguity across all models (14%-20%) suggest that legal terminology disambiguation remains a challenging problem even with our enhanced framework. These findings inform our proposed mitigation strategies and future research directions.
4.2.8 Case Study Analysis: Liebeck v. McDonald’s Hot Coffee Burn Case